OpenStudy (anonymous):

soybean meal is 14% proteincornmeal is 7% proyein. how many pounds of each should be mixed together in order to get 28% mixture that is 8% protein

OpenStudy (anonymous):

what is the meaning

OpenStudy (anonymous):

sorry 280 pounds not 28%

OpenStudy (blacksteel):

Okay, we want to mix together x pounds of soybean meal and y pounds of cornmeal. This gives us x + y = 280. We also know that we want the final mixture to be 8% protein, so we have that 0.14x + 0.07y = 0.08*280. This gives us a system of two equations in 2 unknowns, so it's a simple substitution problem: y = 280 - x. 0.14x + 0.07(280 - x) = 0.08*280 0.14x - 0.07x = 0.08*280 - 0.07*280 0.07x = 0.01*280 x = 40. Then y = 280 - 40 = 240. We can check this easily: 240*0.07 + 40*0.14 = 16.8 + 5.6 = 22.4 = 0.08*280.